OpenStudy (anonymous):

Trying to solve this pls help... Suare root of 11x-3=square root of x +2?? i dnt want just answer i would like to see the steps

OpenStudy (anonymous):

\[\sqrt{11x-3}=\sqrt{x+2}\]

OpenStudy (anonymous):

Is that the problem

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright, well if you square both sides that will get rid of the square root... 11x-3=x+2

OpenStudy (anonymous):

yea i got that step down but then? do i like (11x-3)(11x-3)=(x=2)(x=2)

OpenStudy (anonymous):

****+

OpenStudy (anonymous):

No, you just solve for x add the 3 to both sides 11x=x+5 subtract x 10x=5 divide by 10 x=5/10 x=1/2

OpenStudy (anonymous):

ohhhh okkkkkk thanks so much!!! it was so simple lol i just need to refresh on this for my final

OpenStudy (anonymous):

You're welcome